I need help on this one please.
9. If 2 cards are drawn (without replacement) from a deck of 52 shuffled cards, what is the probability that both cards are kings?

Answer by rothauserc(4718)

There are four kins in the deck of 52
:
Probability that both cards are kings = (4/52) * (3/51) = (1/13) * (1/17) = 1/221 = 0.0045
: